OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = x − 2, find f[g(−6)

OpenStudy (espex):

What is is telling you, is that if you were to replace 'x' with -6, in your g(x) equation, then take that result and plug into f(x), what would you get?

OpenStudy (espex):

So take your g(x), which is x-2, and plug in -6 for 'x' and solve.

OpenStudy (anonymous):

well i already worked out the problem and i got 3 and i wanted to know if that was right?

OpenStudy (espex):

Ah, no, that is not correct. Care to walk me through your steps?

OpenStudy (anonymous):

im not quite sure how i did it. it was mostly mental math. so let me see if i understand. so for the first part i got -8, then do i bring that over to the other side?

OpenStudy (espex):

You take your -8 and plug it in for the 'x' in the expression, x-4

OpenStudy (anonymous):

i just realized that i copied the question wrong! the actual question is: f f(x) = 1/2x − 4 and g(x) = x − 2, find f[g(−6)].

OpenStudy (anonymous):

sorry!!!

OpenStudy (espex):

haha, well that does change things, however 3 is still not a correct answer. :) Take your -8 and plug it in for your 'x' in \(\frac{1}{2}x-4\)

OpenStudy (anonymous):

so -4-4?

OpenStudy (espex):

Exactly. Nice work. :)